OpenStudy (anonymous):

(3a)(2a) and a=-1 help me please

OpenStudy (anonymous):

Put is this way (3 * -1) * (2*-1)

OpenStudy (anonymous):

Now, tell me, what is 3 times -1

OpenStudy (anonymous):

and what is 2 times -1

OpenStudy (anonymous):

once you get the two answers, multiply them together and that's your answer

OpenStudy (anonymous):

just plug in a=-1 and then simply multiply them and remeber.. -1*-1=+1 -1*+1=-1 +1*-1=-1 +1*+1=1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem